OpenStudy (anonymous):

What is the equation of a parabola with vertex (0,0) and directrix x= -3

OpenStudy (anonymous):

y^2=4(3)x

OpenStudy (anonymous):

for y^2=4ax the directrix is given by x=-a

OpenStudy (anonymous):

a proper diagram will be better